OpenStudy (anonymous):

which of the following relations is a function a:(1,4),(1,-6),(1,3),(-7,2) b:(1,4),(-1,2),(5,1),(-7,2) c:(5,1),(-1,4),(1,1),(5,2) D:(1,0)(-1,3),(5,1),(-1,5)

OpenStudy (anonymous):

please someone help?

OpenStudy (ranga):

Do the vertical line test. For a function same x value should NOT give different y values.

OpenStudy (anonymous):

what do you mean may I ask?

OpenStudy (ranga):

Let us look at a:(1,4),(1,-6),(1,3),(-7,2) (1,4) means when x = 1, y = 4 (1,-6) means when x = 1, y = -6 The same x value of 1 gives two different y values. This is NOT a function.

OpenStudy (anonymous):

so the answer would be d or c?

OpenStudy (ranga):

In c: look at the first and last points.

OpenStudy (anonymous):

yeah I see it first point 5,1 and last one 5,2 that means c is the answer?

OpenStudy (ranga):

No. The same x = 5 gives two different y's: 1 and 2. NOT a function.

OpenStudy (ranga):

For a function one x value should give one and only one y value.

OpenStudy (anonymous):

so the answer be d then,soory im kind of new with this type of math.

OpenStudy (ranga):

Look at the second and fourth point in D

OpenStudy (anonymous):

yeah -1,3 and -1,5

OpenStudy (ranga):

same x giving two different y's. NOT a function.

OpenStudy (anonymous):

yeah im starting to see the error

OpenStudy (ranga):

so the answer is b.

OpenStudy (anonymous):

thanks hey just one more question

OpenStudy (anonymous):

simplify the following expression (4^3)-2 a 1/4096 b4 c4,096 d1/4

OpenStudy (anonymous):

I think its option a but correct if im wrong

OpenStudy (ranga):

correct. (4^3)^(-2) = (64)^(-2) = (1/64)^2 = 1/4096